OpenStudy (maddiie):

Find the soultion to each system by the addition(elimination) method. 4s-5t=18 6s + 10t=-8

OpenStudy (watchmath):

Multiply the first equation by 2 and the add it to the second equation to get \(14s=28\) \(s=2\) \(8-5t=18\) \(-5t=10\) \(t=-2\)

OpenStudy (anonymous):

(4s-5t=18)2 6s+10t=-8 ____________ 8s-10t=36 6s+10t=-8 ____________ 14s = 28 s= 2 4(2)-5t=18 8-18=5t -10=5t t= -2